# Yitian Shi, Edgar Welte, Maximilian Gilles, Rania Rayyes

### v2

@@ Abstract
Grasp learning in noisy environments, such as occlusions, sensor noise, and out-of-distribution (OOD) objects, poses significant challenges. Recent learning-based approaches focus primarily on capturing aleatoric uncertainty from inherent data noise. The epistemic uncertainty, which represents the OOD recognition, is often addressed by ensembles with multiple forward paths, limiting real-time application. In this paper, we propose an uncertainty-aware approach for 6-DoF grasp detection using evidential learning to comprehensively capture both uncertainties in real-world robotic grasping. As a key contribution, we introduce vMF-Contact, a novel architecture for learning hierarchical contact grasp representations with probabilistic modeling of directional uncertainty as von Mises–Fisher (vMF) distribution. To achieve this, we derive and analyze the theoretical formulation of the second-order objective on the posterior parametrization, providing formal guarantees for the model’s ability to quantify uncertainty and improve grasp prediction performance. Moreover, we enhance feature expressiveness by applying partial point reconstructions as an auxiliary task, improving the comprehension of uncertainty quantification as well as the generalization to unseen objects. In the real-world experiments, our method demonstrates a significant improvement by 39% in the overall clearance rate compared to the baselines. Video: https://www.youtube.com/watch?v=4aQsrDgdV8Y
@@ III-A Problem formulation
Figure 2: Illustration of posterior update for grasp baseline direction wrt. predictive distributional uncertainty. We denote $R_{\delta_{t}}^{\top}$ as the rotation matrix to transform $\mathbf{b}$ to the $\mathbf{t}$-th bin.
@@
$\mu^{\prime}_{0}=\frac{\kappa_{0}\mu_{0}+m_{\mathbf{c}}\mu_{\mathbf{c}}}{\kappa_{0}+m_{\mathbf{c}}},\kappa^{\prime}_{0}=\kappa_{0}+m_{\mathbf{c}}$
@@ III-C Bayesian loss for vMF posterior
In general, the expected log-likelihood intends to reduce the epistemic uncertainty by maximizing ($\tanh^{-1}(\kappa^{\prime}_{0})-{\kappa^{\prime}_{0}}^{-1}$), while in Bayesian loss, this is balanced through the maximization of the entropy $\mathbb{H}(\mathrm{vMF}(\mu_{\mathbf{c}}|\mu^{\prime}_{0},\kappa^{\prime}_{0}))$ in Eq. (1) (ii) that plays as normalization, which can be derived as [30]:
@@
Figure 3: vMF-Contact architecture. The raw point clouds from a single camera view go through the PointNet-based backbone (I), where the feature is further enhanced by geometric-aware self-attention. The down-scaled point features are taken by (II) pointwise linear layers to predict the conditional grasp orientations. A residual flow is trained to estimate the density of point features (III) and provide the evidence that serves the posterior update (V). Auxiliary point completion (IV) is applied using a shared folding net to enhance the feature expressiveness (green points) and robustness against input noises supervised by ”ground truths” point clouds (blue).
Base on our analysis, vMF-Contact is built upon NatPN and able to enhance feature learning for probabilistic contact grasp through a Bayesian treatment. To ensure seamless integration with point cloud-based learning, our vMF-Contact architecture is designed to incorporate evidential learning to capture the uncertainty under contact grasp representation.
@@
Moreover, recent works [6, 34] have emphasized the importance of enhancing feature understanding in posterior networks through two key strategies: First, the complexity of the encoder architecture and the expressiveness of its features generally lead to improved uncertainty quantification. Based on this, we selected PointNeXt as our backbone following a comprehensive comparison with state-of-the-art PointNet-based encoders in our experiments (Sec. IV). Second, to adapt an analogous auxiliary reconstruction task [6] to the 3D PointNet-based architecture, we employed a shared FoldingNet [35] to reconstruct the local point clouds around the generated contact points, using the feature with feature $z_{\mathbf{c}}$. This aims to extend the pixel-wise reconstruction from 2D features as suggested by [6] to the point-based domain. We supervise the reconstruction results by uniformly sampled points from the object model without occlusion, which aims to improve robustness against input noise while emphasizing local geometric details. It is worth noting that, our reconstruction head is used solely as auxiliary supervision, which is not involved during inference.
@@ III-E3 Flow and reconstruction
Figure 4: i) Generated data from simulation with ground truths point clouds ii) grasps in the scene iii) ID objects iv) OOD objects
@@ IV-C Simulation experiments
Here, we focus on the prediction error and uncertainty estimation of baseline vectors. For baseline vector error, we applied the cosine similarity score as the first metric. Regarding uncertainty calibration error, we utilize the area under the sparsification curve (AUSC) [41] to evaluate the predicted uncertainty, along with the area under the sparsification error (AUSE), which measures the discrepancy between the AUSC and the actual error curve sorted by the defined error metric. Specifically, we sort the predicted contact points based on the uncertainty scores and calculate the error by top $k\%$ samples with $k$-th percentile. Same as the first metric, we take cosine similarity as the error metric for the actual error curve between the predicted contact baseline and its matched ground-truth.
@@ IV-D Realworld experiments
In our real-world experiments, we aim to demonstrate the robustness of our method in handling uncertainty arising from novel scenarios. We utilize a UR10e robotic manipulator equipped with an Orbbec Femto Mega low-cost camera system and a Robotiq 2F85 gripper, as depicted in Fig. 4. In addition, we conduct two sets of experiments: one with in-distribution (ID) (cf. Fig. 4 iii) objects, which are the same objects used for training data generation, and another one with out-of-distribution (OOD) items (cf. Fig. 4 iv). The OOD items differ from the ID objects by various geometrical and physical properties such as a soft cube and a deformable glove.
TABLE II: Real-world experiment results on PointNeXt
Condition Average Success Rate % Average Clearing Rate %
@@
none 39.23 45.45 61.11 66.67
recon 60.61 56.09 83.33 77.78
recon + unc 72.72 65.00 88.89 88.89
@@
During the inference, we take the PointNext-based architecture as the baseline. The generated grasps are filtered using two strategies: i) the grasp quality $\text{P}(\mathbf{c})$ or ii) total uncertainty $\kappa^{\prime}_{0}$. For execution, we randomly select the grasp from $10$-best scored grasps to avoid inference stagnation. We utilize grasp success rates and clearing rates as general performance metrics. in our real-world experiments. Three separate experiments were conducted using the same set of objects and scenes to enable quantitative comparison: 1) baseline with PointNeXt-based architecture, 2) with reconstruction, 3) with reconstructions and uncertainty. In total, we conducted 12 experiments.
@@ V Conclusions
We proposed vMF-Contact, an evidential grasp learning framework that leverages point-wise features to model both aleatoric and epistemic uncertainties with Bayesian treatment. We employed vMF prior to modeling directional uncertainty in the contact grasp representation and formulated a principle objective to implement posterior update. We further enhance feature expressiveness by applying patch-wise point reconstructions, which not only strengthens uncertainty quantification but also significantly enhances the model’s generalization capability to handle OOD objects with varying scenes, viewpoints and sensor noises. Our method is proven to be valid across various PointNet-based architectures according to the evaluation. The real-world experiments further demonstrate the effectiveness of our methods incorporating reconstruction and uncertainty-based sorting significantly improves grasp success.

### v3

@@ Abstract
Grasp learning in noisy environments, such as occlusions, sensor noise, and out-of-distribution (OOD) objects, poses significant challenges. Recent learning-based approaches focus primarily on capturing aleatoric uncertainty from inherent data noise. The epistemic uncertainty, which represents the OOD recognition, is often addressed by ensembles with multiple forward paths, limiting real-time application. In this paper, we propose an uncertainty-aware approach for 6-DoF grasp detection using evidential learning to comprehensively capture both uncertainties in real-world robotic grasping. As a key contribution, we introduce vMF-Contact, a novel architecture for learning hierarchical contact grasp representations with probabilistic modeling of directional uncertainty as von Mises–Fisher (vMF) distribution. To achieve this, we analyze the theoretical formulation of the second-order objective on the posterior parametrization, providing formal guarantees for the model’s ability to quantify uncertainty and improve grasp prediction performance. Moreover, we enhance feature expressiveness by applying partial point reconstructions as an auxiliary task, improving the comprehension of uncertainty quantification as well as the generalization to unseen objects. In the real-world experiments, our method demonstrates a significant improvement by 39% in the overall clearance rate compared to the baselines. The code is available under: https://github.com/YitianShi/vMF-Contact/
@@ III-A Problem formulation
Figure 2: Illustration of posterior update for grasp baseline direction wrt. predictive distributional uncertainty. We denote $R_{t}$ as the rotation matrix to transform $\mathbf{b}$ to the $\mathbf{t}$-th bin.
@@
$\mu^{\prime}_{0}=\frac{\kappa_{0}\mu_{0}+m_{\mathbf{c}}\mu_{\mathbf{c}}}{\kappa_{0}+m_{\mathbf{c}}},\quad\kappa^{\prime}_{0}=\kappa_{0}+m_{\mathbf{c}}$
@@ III-C Bayesian loss for vMF posterior
For the minimization of Eq. 1, this term intends to reduce the epistemic uncertainty (increase ${\kappa^{\prime}_{0}}$) by maximizing ($\tanh^{-1}(\kappa^{\prime}_{0})-{\kappa^{\prime}_{0}}^{-1}$). While this is balanced through the maximization of the term (ii), as the posterior’s entropy $\mathbb{H}(\mathrm{vMF}(\mu_{\mathbf{c}}|\mu^{\prime}_{0},\kappa^{\prime}_{0}))$. This can be derived as [30]:
@@
Figure 3: vMF-Contact architecture. The raw point clouds from a single camera view go through the PointNet-based backbone (I), where the feature is further enhanced by geometric-aware self-attention. The down-scaled point features are taken by (II) pointwise linear layers to predict the conditional grasp orientations. A residual flow is trained to estimate the density of point features (III) and provide the evidence that serves the posterior update (IV). Auxiliary point completion (IV) is applied using a shared folding net to enhance the feature expressiveness (green points) and robustness against input noises supervised by ”ground truths” point clouds (blue).
vMF-Contact is built upon NatPN and able to enhance feature learning for probabilistic contact grasp through a Bayesian treatment. To ensure seamless integration with point cloud-based learning, our vMF-Contact architecture is designed to incorporate evidential learning to capture the uncertainty under contact grasp representation.
@@
Moreover, recent works [6, 34] have emphasized the importance of enhancing feature understanding in posterior networks through two key strategies: First, the complexity of the encoder architecture and the expressiveness of its features generally lead to improved uncertainty quantification. Based on this, we selected PointNeXt as our backbone following a comprehensive comparison with state-of-the-art PointNet-based encoders in our experiments (Sec. IV). Second, to adapt an analogous auxiliary reconstruction task [6] to the 3D PointNet-based architecture, we employed a shared FoldingNet [35] to reconstruct the local point clouds around the generated contact points (V), using the feature with feature $z_{\mathbf{c}}$. This aims to extend the pixel-wise reconstruction from 2D features as suggested by [6] to the point-based domain. We supervise the reconstruction results by uniformly sampled points from the object model without occlusion, which aims to improve robustness against input noise while emphasizing local geometric details. It is worth noting that, our reconstruction head is used solely as auxiliary supervision, which is not involved during inference.
@@ III-E3 Flow and reconstruction
Figure 4: i) Generated data from simulation with ground truths point clouds ii) Ground truths grasps in the scene iii) ID objects iv) OOD objects
@@ IV-C Simulation experiments
Here, we focus on the prediction error and uncertainty estimation of baseline vectors. For baseline vector error, we applied the cosine similarity score as the first metric. Regarding uncertainty calibration error, we utilize the area under the sparsification curve (AUSC) [41] to evaluate the predicted uncertainty, along with the area under the sparsification error (AUSE), which measures the discrepancy between the AUSC and the actual error curve sorted by the defined error metric. Specifically, we sort the predicted contact points based on the uncertainty scores and calculate the error by top $k\%$ samples with $k$-th percentile. As with the first metric, we take cosine similarity as the error metric for the actual error curve between the predicted contact baseline and its matched ground truth.
@@ IV-D Realworld experiments
In our real-world experiments, we aim to demonstrate the robustness of our method in handling uncertainty arising from novel scenarios. We utilize a UR10e robotic manipulator equipped with an Orbbec Femto Mega low-cost camera system and a Robotiq 2F85 gripper, as depicted in Fig. 4. In addition, we conduct two sets of experiments: one with in-distribution (ID) (cf. Fig. 4 iii) objects, which are the same objects used for training data generation, and another one with out-of-distribution (OOD) items (cf. Fig. 4 iv). The OOD items differ from the ID objects by various geometrical and physical properties such as a soft cube, a deformable glove and a cordless screwdriver.
TABLE II: Real-world experiment results
Condition Success Rate % Clearing Rate %
@@
none 39.2 45.5 61.1 66.7
recon 60.6 56.1 83.3 77.8
recon + unc 72.7 65.0 88.9 88.9
@@
During the inference, we take the PointNext-based architecture as the baseline. The generated grasps are filtered using two strategies: i) the grasp quality $\text{P}(\mathbf{c})$ or ii) total uncertainty $\kappa^{\prime}_{0}$. For execution, we randomly select the grasp from $10$-best scored grasps to avoid inference stagnation. We utilize grasp success rates and clearing rates as general performance metrics. in our real-world experiments. Three separate experiments were conducted using the same set of objects and scenes to enable quantitative comparison: 1) baseline with PointNeXt-based architecture, 2) with reconstruction, 3) with reconstructions and uncertainty. In total, we conducted 18 experiments.
@@ V Conclusions
We proposed vMF-Contact, an evidential grasp learning framework that leverages point-wise features to model both aleatoric and epistemic uncertainties with Bayesian treatment. We employed vMF prior to modeling directional uncertainty in the contact grasp representation and formulated a principle objective to implement posterior update. We further enhance feature expressiveness by applying patch-wise point reconstructions, which not only strengthens uncertainty quantification but also significantly enhances the model’s generalization capability to handle OOD objects with varying scenes, viewpoints and sensor noises. Our method is proven to be valid across various PointNet-based architectures according to the evaluation. The real-world experiments further demonstrate the effectiveness of our methods incorporating reconstruction and uncertainty-based sorting, significantly improves grasp success.